OpenStudy (anonymous):

What is the value of 4C2?

OpenStudy (anonymous):

6 8 16 80

OpenStudy (hoblos):

nCr = n!/r!(n-r)! 4C2 = 4!/2!2! = 24/4 = 6

OpenStudy (anonymous):

thanks

OpenStudy (hoblos):

yw =)